The Sad State Lottery requires you to select a sequence of three different numbers from zero through 53. (Order is important.) You are a winner if your sequence agrees with that in the drawing, and you are a booby prize winner if your selection of numbers is correct, but in the wrong order. (Round all answers to three significant figures. Enter the answers in scientific notation.)
What is the probability of being a winner?
1
What is the probability of being a booby prize winner?
2
What is the probability that you are either a winner or a booby prize winner?
3

Answer by edjones(7569)

nPr=n!/(n-r)!
.
1)
nPr=n!/(n-r)!
1/(53!/(53-3)!)=1/140,556
.
2)
nCr=n!/((n-r)!r!)
1/(53!/(50!*3!))-(53!/(53-3)!)
=1/23,426 - 1/140556
=6/140556 - 1/140556
=5/140556
.
3)
5/140556 + 1/140,556
=6/140,556
.
Ed